| 
Group Selection Page
HELP
Controls Group
ITDM
Vacuum
WP78
FXE
SPB
HED
SQS
SXP
Sample Environment
PSPO
XO
TS
Migrated2Zulip
Archived2GitlabPages
Archived2PDF
MigratedByGroup
 | 
| 
SCS
Photon Commissioning Team
XRO
DOC
WP74
MID
XO_mig
XO_Operation
SQS Migrated2Zulip
DET
 | 
| 
Technical integration and coordination
 | 
|   Technical integration and coordination logbook, Contact: A. Violante, G. Wellenreuther  | 
Not logged in | 
  | 
   | 
| 
Message ID: 211
    Entry time: 29 Mar 2021, 10:45
 |  
| Author: | 
Frederik Wolff-Fabris  | 
 | Location: | 
General  | 
 | Severity: | 
Information  | 
 | Subject: | 
Notes XO Integration Meeting 29.03.2021  | 
 
  |  
 
 | 
General Announcement:
- The UOM does not take place on next Friday 02.April.
- Pouneh S. informs the SCS panic button is not properly working (the search hutch does not perform as expected). This does not prevent operation and will be look after 
in the coming days.
CTRL:
- Following the DOC developments; nothing major to be announced.
- Karabo 2.10.4 is to be released which contains features to facilitate opening scenes and has some bug features fixed. CTRL hopes for a wider adoption. 
DA:
- Nothing major to report.
EEE:
- EEE has a time slot on Wednesday from 10:00 to 15:00 to update loops in SA3 and SA2. Updates shall be transparent and an e-mail will be sent. 
A. Galler confirms there will be no beam operation during this time.
VAC:
- Nothing to report.
XPD:
- ZZ request for Shimadzu camera in XTD9: A. Galler says it is not yet scheduled. As soon as valid information is coming, it will be communicated (maybe on Wednesday 
during the loop updates time slot).
DET:
- Small things were fixed with DOC help on available calibration constants for one SPB detector was fixed.
ITDM: 
- Reminder of tomorrow's maintenance on storage system.
XRO:
- Not present.
LAS:
- Thursday's SA2 ADC problem was fixed. Today LAS and EEE will attempt to fix the SA3 missing triggers in the ADC cards.
- All PP-lasers are running.
UNSYS:
- Information: there will be from today to Wednesday tests in SA2 performance with and without air conditioning systems.
SPB:
- Few problems with last week solved by DOC.
- SPB shift crew said machine was running at 2.2MHz. This had to be informed to BKR Operators.
FXE:
- Nothing to report.
MID:
- MID requests that planned interventions in MID hutches that are longer than few minutes and restrict access to it shall be informed in advance. This allows proper 
planning of activities. 
HED:
- Started using PP-laser and preparing for next user run (also tests with THz laser).
- Konstantin S. would like to thank EEE for the help received during last week.
SCS:
- Nothing to report.
SQS:
- Nothing to report.  |